(iv) Meena went to a bank to withdraw Rs.2000. She asked the cashier to give her Rs.50 and Rs.100 notes only. Meena got 25 notes in all. Find how many notes of Rs.50 and Rs.100 she received.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let the number of Rs.50 notes be A and the number of Rs.100 notes be B According to the given information, A + B = 25 ……………………………………………………………………….. (i) 50A + 100B = 2000 ………………………………………………………………(ii) When equation (i) is multiplied with (ii) we get, 50A + 50B = 1250 …………………………………………………………………..(iii) Subtracting the equation (iii) from the equation (ii) we get, 50B = 750 B = 15 Substituting in the equation (i) we get, A = 10 Hence, Meena has 10 notes of Rs.50 and 15 notes of Rs.100.